OpenStudy (anonymous):

draw a graph of the function and indicate the average value \[f(x)=x ^{3}-5x ^{2}+10; \left[ 0,4 \right]\]

OpenStudy (anonymous):

average value is \[\frac{f(4)-f(0)}{4-0}\]

OpenStudy (anonymous):

don't I use the anti-derivative of the function?

OpenStudy (anonymous):

oh i see you want \[\frac{1}{4}\int_0^4(x^3-5x+10)dx\]

OpenStudy (anonymous):

aka "mean value" right

OpenStudy (anonymous):

yes I believe that is what the question calls for

OpenStudy (anonymous):

ok the "anti derivative" is not too bad because this is a polynomial it is \[\frac{x^4}{4}-\frac{5x^2}{2}+10x\]

OpenStudy (anonymous):

(5x^3)/3 right

OpenStudy (anonymous):

plug in 4 ignore 0 because that just gives you 0 then divide by 4

OpenStudy (anonymous):

right i did all that. i got approx. -.75. does that seem correct?

OpenStudy (anonymous):

it is \(-5x\) or \(-5x^2\) ?

OpenStudy (anonymous):

-5x^2

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

then no it is \(-\frac{2}{3}\)

OpenStudy (anonymous):

ok i get that. But how would they go on the graph. Could you give me a crude drawing?

OpenStudy (anonymous):

here is a nice drawing http://www.wolframalpha.com/input/?i=x^3-5x^2%2B10+domain+0..4

OpenStudy (anonymous):

and the average value represented on the graph would just be a straight line at -2/3?

OpenStudy (anonymous):

that i do not know sorry

OpenStudy (anonymous):

ok thank you